Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
Three batteries are connected in series so that the total voltage is 54 volts. The voltage of the first battery is twice the voltage of the second and 1/3 the voltage of the third battery. Find the actual voltage of each battery.

Answer by checkley77(12844)

LET THE 3 BATTERIES BE X, Y & Z
X=2Y GIVEN.
Z=3X OR Z=3(2Y) GIVEN.
Y=Y ASSIGNED.
X+Y+Z=54 GIVEN.
2Y+Y+3(2Y)=54
2Y+Y+6Y=54
9Y=54
Y=54/9
Y=6 VOLTS FOR THE MIDDLE BATTERY
X=2*6=12 VOLTS FOR THE FIRST BATTERY
Z=3*12
Z=36 FOR THE THIRD BATTERY
PROOF:
12+6+36=54
54=54